OpenStudy (anonymous):

{I dont get how to do this, help?} The cylinders are similar. Find the volume of the smaller cylinder. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

I know how to get the volume of a normal shape but this ones weird..

OpenStudy (anonymous):

The complete problem.

OpenStudy (anonymous):

depends what grade/level math this is. chances are it's supposed to be treated as a circle

OpenStudy (anonymous):

around 12th grade math.

OpenStudy (anonymous):

i don't get how to do this since im missing the number to get the volume. And what do i do with the bigger shape?

OpenStudy (anonymous):

i'd say treat as circle V = pi * r^2 * h

OpenStudy (anonymous):

V=3.14*2.9*11.6= Like that?

OpenStudy (anonymous):

and since we know they are similar, that means that their ratios of height to radius (or any other dimensions) are the same. \[\frac{ r_1 }{ h_1 } = \frac{ r_2 }{ h_2 }\]

OpenStudy (anonymous):

you must square the radius: \[V = \pi * r^2 * h\]

OpenStudy (anonymous):

finding height of the smaller one: \[\frac{ 5.8 }{ 11.6 } = \frac{ 2.9 }{ h_2 }\]

OpenStudy (anonymous):

whats the h2 ?

OpenStudy (anonymous):

height of cylinder 2

OpenStudy (anonymous):

r is radius